OpenStudy (anonymous):

Evaluate the expression Arcsec(sqrt 2). Do not use your calculator to answer this question. Do not express answer in decimals.

OpenStudy (anonymous):

arcsec(sqrt2)=x sec(x) = sqrt(2) 1/cos(x) = sqrt(2) 1/sqrt(2) = cos(x) using the special triangle with side lengths 1,1 and sqrt(2), and angles of pi/4, we can see that x= pi/4. therefore arcsec(sqrt(2)) = pi/4.

OpenStudy (anonymous):

thanks